# USTC-ICTS/PCFT-24-24

### v1

@@ I Introduction
Boson stars are nowadays well known solutions that arise in the context of complex scalar fields coupled to Einstein gravity. The first such self-gravitating, spatially localized, stationary configurations were obtained more than 50 years ago Kaup (1968). This particular nonlinear solution, which nowadays is known as a mini-boson star, arise within the simplest model of a minimally coupled free scalar field and when spherical symmetry is assumed. Self-interactions such as the quartic Colpi et al. (1986) and sextic Friedberg et al. (1987) terms in the scalar potential enriches the family of boson stars. The variety of boson stars that can be found in the literature goes beyond modifications to the scalar potential, see e.g. Jetzer and van der Bij (1989) for a generalization including electrically charged scalar fields, Herdeiro and Radu (2014) for solutions with horizons, Brito et al. (2016) for solutions of a vectorial bosonic field and some others to which we will refer later. However, many of them (and in particular the self-interacting ones) share similar dynamical properties with the mini-boson stars (see Liebling and Palenzuela (2023) for a review on this topic), the most important of which include the stability and existence of a formation mechanism Seidel and Suen (1990); Balakrishna et al. (1998); Seidel and Suen (1994); Bezares et al. (2017); Sanchis-Gual et al. (2022); Jaramillo et al. (2020); Sanchis-Gual et al. (2021); Bezares and Palenzuela (2018); Croft et al. (2023); Atteneder et al. (2024); Sanchis-Gual et al. (2019a); Helfer et al. (2022); Palenzuela et al. (2007); Gleiser and Watkins (1989); Cunha et al. (2017); Kleihaus et al. (2012); Tamaki and Sakai (2010). This has enable boson stars to find applications in various astrophysical and cosmological scenarios Olivares et al. (2020); Amaro-Seoane et al. (2010); Bernal and Siddhartha Guzman (2006); Grandclement et al. (2014); Suárez et al. (2014); Hui et al. (2017); Jaramillo and Núñez (2022); Choi et al. (2019); Solís-López et al. (2021); Annulli et al. (2020); Guzman and Urena-Lopez (2006); Guzman and Rueda-Becerril (2009); Bernal and Matos (2024)
@@
Q-balls Friedberg et al. (1976); Coleman (1985) (see Lee and Pang (1992) for a review) are the Minkowski space cousins of boson stars. Recently, it has been observed that for the model where spherical Q-balls exist, there also exist a tower of complex, multipolar Q-balls Copeland et al. (2014). These “molecular states” of Q-balls, despite being metastable, can live for a long time Copeland et al. (2014); Xie et al. (2021), especially for the logarithmic potential Hou et al. (2022). These complex Q-balls have the intriguing feature that within a ball the co-existing positive and negative charges swap with time. The charge-swapping Q-balls are attractor solutions that can form from quite generic initial conditions Xie et al. (2021); Hou et al. (2022), and in particular they may form as a consequence of the Affleck-Dine condensate fragmentation Hou et al. (2022); Zhou (2015), thanks to parametric resonance.
@@
We also find that, while it is possible to achieve a considerable compactness, those charge-swapping boson stars in the large-size, low-compactness limit are well suited for a candidate for galactic dark matter halos. In such a scenario, the morphology of the dipolar configuration influences the structures moving in the outer region of the halo, and as a result, the galactic satellites moving around a host galaxy travel in coherently oriented planes. We shall explore the possibility of explaining the anisotropic distribution of satellite galaxies Pawlowski (2021) by a charge-swapping dipolar galactic halo. These charge-swapping structures can be formed as a result of collisions of two structures with opposite charges, or, as mentioned, they can be formed by fragmentation of Affleck-Dine-like condensate in earlier universe.
@@ II Theory and setup
where the Newton’s constant $G=1/m_{P}^{2}$ and
@@
[]delimited-[][$\dots$][ … ] in the equation for $\tilde{A}_{ij}$ and $\tilde{\Gamma}^{i}$ denote the standard quantities defined in the right-hand side of the BSSNOK equations, except for the source terms which have been explicitly included above. The real and imaginary part of the scalar field $\phi$ are evolved separately in the actual code. Finally, the evolution is subject to a set of constraints given by
@@
II.2 Stationary isolated Boson Stars
@@
For every pair of $(\alpha_{G},\mathfrak{g})$, there exists a family of solutions parameterized (in most cases) by the value of the scalar field at the center of the star, $f_{0}$. For $\alpha_{G}\ll 1$, we expect to recover the Q-ball solutions, and, in particular, the parameter space explored in Xie et al. (2021). In Fig. 1, we plot the mass and the compactness for sequences of solutions with different values of coupling constant $\alpha_{G}$. From the left panel of Fig. 1 we see that, contrary to what happens in the Q-ball case, the mass of the solutions is regularized at the maximal frequency $\omega$. Additionally, when gravity is introduced, solutions beyond the Q-ball thin-wall limit can exist. In our case, this limit is $\omega=(1-1/(4\mathfrak{g}))^{1/2}$. These solutions are of particular interest since, as can be seen in the right panel of Fig. 1, the compactness can reach very high values. For the corresponding static black hole, we have $\mathcal{C}=0.5$, while the maximum of $\mathcal{C}$ is around $0.1$ for mini-boson stars ($\lambda=0=\tilde{g}$) Jaramillo et al. (2023) and around $0.16$ for strongly self-interacting boson stars (large $\lambda/(G\mu^{2})$, $\tilde{g}=0$) Amaro-Seoane et al. (2010); Colpi et al. (1986).
@@ II.4 Preparing complex boson stars: solving initial constraints
General Relativity is highly nonlinear, and the gravitational field in general does not obey a linear superposition law. Thus, we can not simply superpose boson star solutions, which is in general not a solution, and let it relax towards a complex boson star configuration. We need to carefully prepare the initial data by solving the Hamiltonian and momentum constraints. The 3+1 formalism with a scalar field as a source was presented in Section II.1. For our purposes, we shall choose $K_{ij}=0$ at $t=0$, which means that the momentum density $P^{i}$ of the initial data that we intend to construct vanishes, that is, we wish to start with two clumps standing initially with zero velocity. The extrinsic curvature vanishes for time symmetric data, and therefore we can ignore the momentum constraints (25) and solve only the Hamiltonian constraint (24), which in this case is reduced to
@@
where $K_{\phi}^{\rm(BS)}=i\omega f^{\rm(BS)}e^{-i\omega t}/(2\alpha^{\rm(BS)})$ and $K_{\phi}^{\rm(aBS)}=-i\omega f^{\rm(BS)}e^{i\omega t}/(2\alpha^{\rm(BS)})$. Then, we make a simple superposition of the corresponding gravitational fields. This setup gives us the initial guess in the Newton iterative scheme used to solve the Hamiltonian constraint.
More concretely, according to Eq. (27), the spatial metric of an isolated boson star in these coordinates is simply $\gamma_{ij}^{(\mathrm{BS})}=\Psi_{(\mathrm{BS})}^{4}\delta_{ij}$. Recall that the background metrics for an isolated boson star and anti-boson star are the same, so a simple superposition of the spatial metric $\tilde{\gamma}_{ij}$ takes the form Helfer et al. (2022); Atteneder et al. (2024),
$\begin{split}\tilde{\gamma}_{ij}(\mathbf{r})&=(\Psi_{(\mathrm{BS})}^{4}(x,y,z-d/2)\\ &\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ +\Psi_{(\mathrm{BS})}^{4}(x,y,z+d/2)-1)\,\delta_{ij}\end{split}$ (41)
We shall choose to solve Eq. (38) using the conformal thin-sandwich approach, where the 3+1 spatial metric is related to the conformal metric $\tilde{\gamma}_{ij}$ by $\gamma_{ij}=\psi^{4}\tilde{\gamma}_{ij}$. So, using spherical coordinates and defining111 Although in the dipolar case both $\phi$ and $K_{\phi}$ depend only on $r$ and $\theta$, we here also keep track of the $\varphi$ angle dependence in the geometric and scalar field quantities so that the generalization to the multipolar case is straightforward to obtain from the equations presented here. For the dipole solutions, both the simple superposition guess $\tilde{\gamma}_{ij}$ and the expected solution $\gamma_{ij}$ for the Hamiltonian constraint are axisymmetric, so we also have $A=A(r,\theta)$.
$\begin{split}A(r,\theta,\varphi)=&\left(\Psi_{(\mathrm{BS})}^{4}(x,y,z-d/2)\right.\\ &\left.+\Psi_{(\mathrm{BS})}^{4}(x,y,z+d/2)-1\right)^{1/2}\,;\end{split}$ (42)
the line element reads,
@@
In this formulation, the Hamiltonian constraint becomes $\tilde{D}^{2}\psi-1/8\psi{}{}^{(3)}\!\tilde{R}=-2\pi\psi^{5}\rho$, where $\tilde{D}$ is the covariant derivative associated with $\tilde{\gamma}$ and ${}^{(3)}\!\tilde{R}$ is the 3-dimensional Ricci scalar of $\tilde{\gamma}_{ij}$. For the initial guess for the conformal factor, we choose $\psi(r,\theta,\varphi)=1$, because of the conformal metric we have chosen. Furthermore, we take as initial data for the lapse function $\alpha$ a simple superposition Palenzuela et al. (2007); Bezares and Palenzuela (2018); Sanchis-Gual et al. (2019a); Jaramillo et al. (2022) $\alpha(r,\theta)=\alpha^{(\mathrm{BS})}(x,y,z-d/2)+\alpha^{(\mathrm{BS})}(x,y,z+d/2)-1$. See Refs. Helfer et al. (2022); Atteneder et al. (2024) for other possibilities.
@@
Notice that it depends on $\psi$ so $\rho$ should be updated at every iteration in the root-finding procedure used in the elliptic solver for the Hamiltonian constraint.
Finally, combining Eqs. (44), (45) and (48), we arrive at the following partial differential equation:
@@
We solve this elliptic partial differential equation using the Kadath library Grandclement (2009, 2010). This spectral solver has been utilized in various theoretical physics scenarios, particularly for constructing axisymmetric boson stars Grandclement et al. (2014); Jaramillo and Núñez (2022). Besides the axisymmetric dipolar case, we will also be interested in solving Eq. (49) without axisymmetry for the quadrupole and octupole cases. To the end, we employ the spherical space facility implemented in the library, which uses spherical coordinates and decomposes the space into spherical shells with the possibility of placing as the last shell one with a compactified radial coordinate such that boundary conditions at infinity can be imposed exactly. Full details of this space can be found Ref. Grandclement (2010) where the Kadath solver was presented.
@@ III Free scalar
Figure 4: Evolution of charge density $-n_{\mu}j^{\mu}$ in a half charge-swapping period plotted in the $xz$ plane for a dipole configuration (top panel). The color scheme uses blue for positive charge, red for negative charge and black for zero charge, with the numerical values ranging ±$\pm 3.6\times 10^{-4}$. We also show the evolution of the total Noether charge in the $z>0$ half space, $Q^{\rm up}$, for the configuration initially constructed by superposing two boson stars with $\omega=\pm 0.98$ and separated by $d=24$ initially (bottom panel). The points in the inset show the five snapshots of the top panel. $T_{0}=2\pi$ is the characteristic time of the scalar field and also roughly the oscillation frequency of the two initial constituent stars. However, the charge-swapping frequency is much larger. The height of the boxes in the top panel snapshots is $\Delta z=28$. The inset can be seen in even greater detail in Fig. 5. In accordance with footnote 2 we use $\alpha_{G}$ to reconstruct the finite and physical quantity associated to the Noether charge.
@@
This can be formally obtained by taking the limit $\alpha_{G}\to\infty$ and $\mathfrak{g}/\alpha_{G}^{2}\to 0$ from the sextic potential. The gravitational attraction makes it possible to have charge-swapping solitons even for this quadratic/free potential. In this model, stable (mini-)boson stars exist from the zero mass limit, $\omega=1$, up to the critical mass Gleiser and Watkins (1989) located at $\omega=0.853$ and222 In the limit $\alpha_{G}\to\infty$ the dimensionless mass $M$, energy $E$ and other quantities scale as $\alpha_{G}^{-1}$, so we include $\alpha_{G}$ factors to make them finite. However, it is important to say that after restoring units with the formulas given in Section II, the physical mass and Noether charge of the solutions in the $\alpha_{G}\to\infty$ limit correspond to the expected mini-boson star finite quantities (as they should be). For instance for the critical mass $\tilde{M}=0.633/(G\mu)$, as expected for the Kaup limit solution. $\alpha_{G}M=0.633$; see Fig. 3. This stability range of boson stars restricts the domain of configurations we can use to prepare “molecular states” of boson stars.
@@
Table 1: Classical observables of the mini-boson stars used in the preparation of charge-swapping boson stars. All of them correspond to low-compactness stars. The value of $E$ is calculated using the integral (14), despite that it only characterizes the total energy of the solution in the weak-field limit. The value $d_{\rm min}$ in the last column corresponds to the minimal distance between the centers of the boson stars in the dipolar case such that a long-living boson-anti boson star configuration can form.
To determine how the system evolves, the initial data are exported into the evolution scheme Eqs. (22), implemented in the open source Einstein Toolkit infrastructure Ein (2023); Loffler et al. (2012) using the McLachan Brown et al. (2009) thorn to evolve the metric quantities and a modified version of the Scalar thorn Cunha et al. (2017); Witek et al. (2021) to evolve the scalar field. To track the formation of apparent horizons we use the AHFinderDirect Thornburg (2004). The evolution is implemented with a grid spacing of $\Delta x^{i}=1.6$ and a Courant factor of $\Delta t/\Delta x=0.125$ and with different fixed refinement levels included in the Carpet thorn Schnetter et al. (2004). In all cases, the size of the computational domain box is at least twice that of $d+2R_{99}$ (the approximate size of the configuration at $t=0$) for the configuration with highest compactness and largest $d$.
The particular size of the domain and the distribution of the refinement level will be specified in the following and kept fixed for all the simulations in the same section for an easy comparison. As discussed in Xie et al. (2021) and Hou et al. (2022), boundary conditions are important when investigating the lifetime of a charge-swapping configuration. In line with this, we have also employed absorbing (radiative) boundary conditions that are usual for boson star simulations. For the scalar field, we use the standard Sommerfield boundary conditions but additionally include the first order correction to the massless scalar field dispersion relation Seidel and Suen (1990): $\partial_{t}K_{\phi}=-\partial_{r}K_{\phi}-K_{\phi}/r+\mu^{2}\alpha\phi/4$ at $r\to\infty$. As for the gauge freedoms, we impose the “Gamma-driver” condition for shift vector $\beta^{i}$ and the “1+log” condition for the lapse $\alpha$.
The interpolation from the initial data spectral coefficients and the analysis of various quantities during the evolution are done using separate thorns also within the Cactus framework. The choices of the numerical setup employed in this work have been tested in Appendix A.

### v2

@@ I Introduction
Boson stars are nowadays well known solutions that arise in the context of complex scalar fields coupled to Einstein gravity. The first such self-gravitating, spatially localized, stationary configurations were obtained more than 50 years ago Kaup (1968). This particular nonlinear solution, which nowadays is known as a mini-boson star, arise within the simplest model of a minimally coupled free scalar field and when spherical symmetry is assumed. Self-interactions such as the quartic Colpi et al. (1986) and sextic Friedberg et al. (1987) terms in the scalar potential enrich the family of boson stars. The variety of boson stars that can be found in the literature goes beyond modifications to the scalar potential, see e.g. Jetzer and van der Bij (1989) for a generalization including electrically charged scalar fields, Herdeiro and Radu (2014) for solutions with horizons, Brito et al. (2016) for solutions of a vectorial bosonic field and some others to which we will refer later. However, many of them (and in particular the self-interacting ones) share similar dynamical properties with the mini-boson stars (see Liebling and Palenzuela (2023) for a review on this topic and Zhang (2024) for an unified view in the nonrelativistic limit), the most important of which include the stability and existence of a formation mechanism Seidel and Suen (1990); Balakrishna et al. (1998); Seidel and Suen (1994); Bezares et al. (2017); Sanchis-Gual et al. (2022); Jaramillo et al. (2020); Sanchis-Gual et al. (2021); Bezares and Palenzuela (2018); Croft et al. (2023); Atteneder et al. (2024); Sanchis-Gual et al. (2019a); Helfer et al. (2022); Palenzuela et al. (2007); Gleiser and Watkins (1989); Cunha et al. (2017); Kleihaus et al. (2012); Tamaki and Sakai (2010); Li et al. (2020). This has enabled boson stars to find applications in various astrophysical and cosmological scenarios Olivares et al. (2020); Amaro-Seoane et al. (2010); Bernal and Siddhartha Guzman (2006); Grandclement et al. (2014); Suárez et al. (2014); Hui et al. (2017); Jaramillo and Núñez (2022); Choi et al. (2019); Solís-López et al. (2021); Annulli et al. (2020); Guzman and Urena-Lopez (2006); Guzman and Rueda-Becerril (2009); Zeng et al. (2024); Rosa and Rubiera-Garcia (2022); Rosa et al. (2023); Bernal and Matos (2024)
@@
Q-balls Friedberg et al. (1976); Coleman (1985) (see Lee and Pang (1992) for a review) are the Minkowski space cousins of boson stars. Recently, it has been observed that for the model where spherical Q-balls exist, there also exist a tower of complex, multipolar Q-balls Copeland et al. (2014). These “molecular states” of Q-balls, despite being metastable, can live for a long time Copeland et al. (2014); Xie et al. (2021, 2024), especially for the logarithmic potential Hou et al. (2022). These complex Q-balls have the intriguing feature that within a ball the co-existing positive and negative charges swap with time. The charge-swapping Q-balls are attractor solutions that can form from quite generic initial conditions Xie et al. (2021); Hou et al. (2022), and in particular they may form as a consequence of the Affleck-Dine(-like) condensate fragmentation in the early universe Hou et al. (2022); Zhou (2015), thanks to parametric resonance. On a slightly different note, it was recently observed that Q-balls and boson stars can induce superradiant amplification for scattering waves even in the spherically symmetrical cases Saffin et al. (2023); Cardoso et al. (2023); Gao et al. (2024).
@@
We also find that, while it is possible to achieve a considerable compactness, those charge-swapping boson stars in the large-size, low-compactness limit are well suited for a candidate for galactic dark matter halos. In such a scenario, the morphology of the dipolar configuration influences the structures moving in the outer region of the halo, and as a result, the galactic satellites moving around a host galaxy travel in coherently oriented planes. We shall explore the possibility of explaining the anisotropic distribution of satellite galaxies Pawlowski (2021) by a charge-swapping dipolar galactic halo. These charge-swapping structures can be formed as a result of collisions of two structures with opposite charges, or, as mentioned, they can be formed by fragmentation of an Affleck-Dine-like condensate in the early universe.
@@ II Theory and setup
where the Newton’s constant is $G=1/m_{P}^{2}$ and
@@
[]delimited-[][$\dots$][ … ] in the equation for $\tilde{A}_{ij}$ and $\tilde{\Gamma}^{i}$ denote the standard quantities defined in the right-hand side of the BSSNOK equations, except for the source terms which have been explicitly included above. The real and imaginary parts of the scalar field $\phi$ are evolved separately in the actual code. Finally, the evolution is subject to a set of constraints given by
@@
II.2 Stationary isolated boson btars
@@
For every pair of $(\alpha_{G},\mathfrak{g})$, there exists a family of solutions parametrized (in most cases) by the value of the scalar field at the center of the star, $f_{0}$. For $\alpha_{G}\ll 1$, we expect to recover the Q-ball solutions, and, in particular, the parameter space explored in Xie et al. (2021). In Fig. 1, we plot the mass and the compactness for sequences of solutions with different values of coupling constant $\alpha_{G}$. From the left panel of Fig. 1 we see that, contrary to what happens in the Q-ball case, the mass of the solutions is regularized at the maximal frequency $\omega$. Additionally, when gravity is introduced, solutions beyond the Q-ball thin-wall limit can exist. In our case, this limit is $\omega=(1-1/(4\mathfrak{g}))^{1/2}$. These solutions are of particular interest since, as can be seen in the right panel of Fig. 1, the compactness can reach very high values. For the corresponding static black hole, we have $\mathcal{C}=0.5$, while the maximum of $\mathcal{C}$ is around $0.1$ for mini-boson stars ($\lambda=0=\tilde{g}$) Jaramillo et al. (2023) and around $0.16$ for strongly self-interacting boson stars (large $\lambda/(G\mu^{2})$, $\tilde{g}=0$) Amaro-Seoane et al. (2010); Colpi et al. (1986).
@@ II.4 Preparing complex boson stars: solving initial constraints
General Relativity is highly nonlinear, and the gravitational field in general does not obey a linear superposition law. Thus, we can not simply superpose boson star solutions, which is in general not a solution, and let it relax toward a complex boson star configuration. We need to carefully prepare the initial data by solving the Hamiltonian and momentum constraints. The 3+1 formalism with a scalar field as a source was presented in Section II.1. For our purposes, we shall choose $K_{ij}=0$ at $t=0$, which means that the momentum density $P^{i}$ of the initial data that we intend to construct vanishes, that is, we wish to start with two clumps standing initially with zero velocity. The extrinsic curvature vanishes for time symmetric data, and therefore we can ignore the momentum constraints (25) and solve only the Hamiltonian constraint (24), which in this case is reduced to
@@
where $K_{\phi}^{\rm(BS)}=i\omega f^{\rm(BS)}e^{-i\omega t}/(2\alpha^{\rm(BS)})$ and $K_{\phi}^{\rm(aBS)}=-i\omega f^{\rm(BS)}e^{i\omega t}/(2\alpha^{\rm(BS)})$. Then, we make a simple superposition of the corresponding gravitational fields. This setup gives us the initial guess in the Newton iterative scheme used to solve the Hamiltonian constraint. In the following we explain the procedure we follow to choose the initial guess for the spacetime metric and then obtain a solution to the Hamiltonian constraint.
We shall choose to solve Eq. (38) using the conformal thin-sandwich approach, where the 3+1 spatial metric is related to the conformal metric $\bar{\gamma}_{ij}$ by $\gamma_{ij}=\psi^{4}\bar{\gamma}_{ij}$. Now, according to Eq. (27), the spatial metric of an isolated boson star in these coordinates is simply $\gamma_{ij}^{(\mathrm{BS})}=\Psi_{(\mathrm{BS})}^{4}\delta_{ij}$. Recall that the background metrics for an isolated boson star and anti-boson star are the same, so we choose the conformal metric $\bar{\gamma}_{ij}$ to be a simple superposition of the spatial metric Helfer et al. (2022); Atteneder et al. (2024),
$\begin{split}\bar{\gamma}_{ij}(\mathbf{r})&=A(\mathbf{r})^{2}\,\delta_{ij}\end{split}$ (41)
where we define the function $A$ in spherical coordinates111 Although in the dipolar case both $\phi$ and $K_{\phi}$ depend only on $r$ and $\theta$, we here also keep track of the $\varphi$ angle dependence in the geometric and scalar field quantities so that the generalization to the multipolar case is straightforward to obtain from the equations presented here. For the dipole solutions, both the simple superposition guess $\bar{\gamma}_{ij}$ and the expected solution $\gamma_{ij}$ for the Hamiltonian constraint are axisymmetric, so we also have $A=A(r,\theta)$. as
$\begin{split}A(r,\theta,\varphi)=&\left(\Psi_{(\mathrm{BS})}^{4}(x,y,z-d/2)\right.\\ &\left.+\Psi_{(\mathrm{BS})}^{4}(x,y,z+d/2)-1\right)^{1/2}\,.\end{split}$ (42)
The full spatial line element in spherical coordinates reads
@@
In the conformal thin-sandwich formulation, the Hamiltonian constraint becomes the Lichnerowicz equation
$\tilde{D}^{2}\psi-1/8\psi{}{}^{(3)}\!\tilde{R}=-2\pi\psi^{5}\rho\,,$ (44)
where $\tilde{D}$ is the covariant derivative associated with $\bar{\gamma}$ and ${}^{(3)}\!\tilde{R}$ is the 3-dimensional Ricci scalar of $\bar{\gamma}_{ij}$. For the initial guess for the conformal factor, we choose $\psi(r,\theta,\varphi)=1$, because of the conformal metric we have chosen. Furthermore, we take as initial data for the lapse function $\alpha$ a simple superposition Palenzuela et al. (2007); Bezares and Palenzuela (2018); Sanchis-Gual et al. (2019a); Jaramillo et al. (2022) $\alpha(r,\theta)=\alpha^{(\mathrm{BS})}(x,y,z-d/2)+\alpha^{(\mathrm{BS})}(x,y,z+d/2)-1$. See Refs. Helfer et al. (2022); Atteneder et al. (2024); Siemonsen and East (2023); Evstafyeva et al. (2023) for other possibilities.
@@
It is important to note that $\rho$ depends on $\psi$, so it must be updated at each iteration in the root-finding procedure used by the elliptic solver for the Hamiltonian constraint. This approach differs from the one typically employed in the context of fluid stars, where energy and momentum are fixed quantities, allowing for straightforward reconstruction of fluid variables. In our case, if $\rho$ is fixed, the scalar field variables can only be reconstructed after solving the Hamiltonian, which involves finding the scalar field that corresponds to the given $\rho$ through a differential equation. This process is complex and can result in initial data that do not accurately represent the physical situation of interest. Specifically, in charge-swapping configurations, we fix the scalar field initial data (39), (40) and treat them as the essential quantities we can control to achieve the desired dynamical configurations. As a consequence, $\rho$ becomes a derived quantity, which can lead to issues of non-uniqueness in the solutions of Eq. (44).
A comprehensive discussion of this issue in the context of initial data construction for boson stars can be found in Ref.Siemonsen and East (2023). In our implementation, we do not encounter alternative solutions and conduct three tests to assess the robustness of the code. First, we observe that the solutions for isolated boson stars ($\psi=1$) act as local attractors in the solution space, given the method used to solve the Lichnerowicz equation (44). Second, we find that as the stars are moved farther apart, the solution converges to the $\psi=1$ case as the distance $d$ increases. Finally, we experiment with different implementations using superpositions other than those in Eq.(42), yet the (physical) metric coefficient $\psi^{4}A^{2}$ consistently converges to the same solution.
In the following we describe in more detail how we solve the Lichnerowicz equation (44) in practice. Combining Eqs. (45), (46) and (49), we arrive at the following partial differential equation:
@@
We solve this elliptic partial differential equation using the Kadath library Grandclement (2009, 2010). This spectral solver has been utilized in various theoretical physics scenarios, particularly for constructing axisymmetric boson stars Grandclement et al. (2014); Jaramillo and Núñez (2022). Besides the axisymmetric dipolar case, we will also be interested in solving Eq. (50) without axisymmetry for the quadrupole and octupole cases. To this end, we employ the spherical space facility implemented in the library, which uses spherical coordinates and decomposes the space into spherical shells with the possibility of placing as the last shell one with a compactified radial coordinate such that boundary conditions at infinity can be imposed exactly. Full details of this space can be found Ref. Grandclement (2010) where the Kadath solver was presented.
@@ III Free scalar
Figure 4: Evolution of charge density $-n_{\mu}j^{\mu}$ in a half charge-swapping period plotted in the $xz$ plane for a dipole configuration (top panel). The color scheme uses blue for positive charge, red for negative charge and black for zero charge, with the numerical values ranging ±$\pm 3.6\times 10^{-4}$. We also show the evolution of the total Noether charge in the $z>0$ half space, $Q^{\rm up}$, for the configuration initially constructed by superposing two boson stars with $\omega=\pm 0.98$ and separated by $d=24$ initially (bottom panel). The inset in the left shows the initial relaxation process. The points in the inset of the right show the five snapshots of the top panel. $T_{0}=2\pi$ is the characteristic time of the scalar field and also roughly the oscillation frequency of the two initial constituent stars. However, the charge-swapping frequency is much larger. The height of the boxes in the top panel snapshots is $\Delta z=28$. The inset can be seen in even greater detail in Fig. 5. In accordance with footnote 2 we use $\alpha_{G}$ to reconstruct the finite and physical quantity associated to the Noether charge.
@@
This can be formally obtained by taking the limit $\alpha_{G}\to\infty$ and $\mathfrak{g}/\alpha_{G}^{2}\to 0$ from the sextic potential. The gravitational attraction makes it possible to have charge-swapping solitons even for this quadratic/free potential. In this model, stable (mini-)boson stars exist from the zero mass limit, $\omega=1$, up to the critical mass Gleiser and Watkins (1989) located at $\omega=0.853$ and222 In the limit $\alpha_{G}\to\infty$ the dimensionless mass $M$, energy $E$ and other quantities scale as $\alpha_{G}^{-1}$, so we include $\alpha_{G}$ factors to make them finite. However, it is important to say that after restoring units with the formulas given in Section II, the physical mass and Noether charge of the solutions in the $\alpha_{G}\to\infty$ limit correspond to the expected mini-boson star finite quantities (as they should). For instance for the critical mass $\tilde{M}=0.633/(G\mu)$, as expected for the Kaup limit solution. $\alpha_{G}M=0.633$; see Fig. 3. This stability range of boson stars restricts the domain of configurations we can use to prepare “molecular states” of boson stars.
@@
Table 1: Classical observables of the mini-boson stars used in the preparation of charge-swapping boson stars. All of them correspond to low-compactness stars. The value of $E$ is calculated using the integral (14), despite that it only characterizes the total energy of the solution in the weak-field limit. The value $d_{\rm min}$ in the last column corresponds to the minimal distance between the centers of the boson stars in the dipolar case such that a long-living boson-anti boson star configuration can form The configurations with “-” in the last column, represent cases not explored or discussed in the dipolar scenario.
To determine how the system evolves, the initial data are exported into the evolution scheme Eqs. (22), implemented in the open source Einstein Toolkit infrastructure Ein (2023); Loffler et al. (2012) using the McLachan Brown et al. (2009) thorn to evolve the metric quantities and a modified version of the Scalar thorn Cunha et al. (2017); Witek et al. (2021) to evolve the scalar field. To track the formation of apparent horizons we use the AHFinderDirect Thornburg (2004). The evolution is implemented with a grid spacing of $\Delta x^{i}=1.6$ and a Courant factor of $\Delta t/\Delta x=0.125$ and with different fixed refinement levels included in the Carpet thorn Schnetter et al. (2004). In all cases, the size of the computational domain box is at least twice that of $d+2R_{99}$ (the approximate size of the configuration at $t=0$) for the configuration with the highest compactness and largest $d$.
The particular size of the domain and the distribution of the refinement level will be specified in the following and kept fixed for all the simulations in the same section for an easy comparison. As discussed in Xie et al. (2021) and Hou et al. (2022), boundary conditions are important when investigating the lifetime of a charge-swapping configuration. In line with this, we have also employed absorbing (radiative) boundary conditions that are usual for boson star simulations. For the scalar field, we use the standard Sommerfield boundary conditions but additionally include the first order correction to the massless scalar field dispersion relation Seidel and Suen (1990): $\partial_{t}K_{\phi}=-\partial_{r}K_{\phi}-K_{\phi}/r+\mu^{2}\alpha\phi/4$ at $r\to\infty$. As for the gauge freedoms, we impose the “Gamma-driver” condition Alcubierre et al. (2003) for the shift vector $\beta^{i}$ and the “1+log” condition Bona et al. (1995) for the lapse $\alpha$.
The interpolation from the initial data spectral coefficients and the analysis of various quantities during the evolution are done using separate thorns also within the Cactus framework. The choices of the numerical setup employed in this work have been tested in the Appendix.